OpenStudy (anonymous):

what percent of 125 is 43.8?

Parth (parthkohli):

Let's say that it is x% 125 * x/100 = 43.8

OpenStudy (anonymous):

so should i multiply 43.8*125/100?

Parth (parthkohli):

Start by dividing both sides by 125, then multiply both sides by 100.

OpenStudy (anonymous):

43.8/125= 0.3504*100=35.04

OpenStudy (anonymous):

what did i do wrong?

OpenStudy (callisto):

:O But... 125 x 35.04% = 43.8

OpenStudy (callisto):

@ParthKohli Hey! You've typed the numbers in wrong order! It's 4380, not 4830!

Parth (parthkohli):

Oh oh sorry

Parth (parthkohli):

<--- confused owl

OpenStudy (anonymous):

so was my answer correct?

Parth (parthkohli):

Yepyep

OpenStudy (anonymous):

thanks!